Joseph combined 7/4c of pancake mix and 3/4c of water to make pancake batter. If he uses 1/4 of a cup to make each pancake, how many pancakes can Joseph make?
Answer by Theo(13342)

he uses 7/4 cups of pancake mix and 3/4 cups of water to make the pancake batter.

total ingredients = 7/4 + 3/4 = 10/4 cups.

if it takes 1/4 of a cup to make each pancake, then (10/4) / (1/4) = 10/4 * 4 = 10 pancakes

he should be able to make 10 pancakes with the batter.